What is the kinetic energy in kJ/mol, of a hypothetical gas with molecular weight of 20.0 g/mol at a temperature of 65.6 C? Do not try to identify the gas.
1
Expert's answer
2019-10-02T08:03:29-0400

The Average kinetic energy per mole of gas can be calculated by using the formula

E = 3/2 RT

Here R= gas constant = 8.314 J mol-1 K-1

T = 65.6+273 K = 338.6K

E = 3/2 * 8.314 * 338.6 = J mol-1

E= 4222.68 J mol-1 = 4.22 kJ mol-1